OpenStudy (anonymous):

k(2,3) f(4,4) use the distance formula.

OpenStudy (anonymous):

do you know the distance formula?

OpenStudy (anonymous):

nope

OpenStudy (amistre64):

do you know the Pythagorean theorum?

OpenStudy (anonymous):

distance not slope

OpenStudy (anonymous):

\[\sqrt{\left(x_1-x\right){}^2+\left(y_1-y\right){}^2}\]

OpenStudy (amistre64):

distance is "length"; and if you know the length of 2 sides of a right angled triangle, you can find the distance .. i mean length ... of the other side

OpenStudy (amistre64):

x moves for 2 to 4; thats a length of 2 y moves from 3 to 4; thats a length of 1 how do we find the length from this?